destinyobioha:

rearrange the formula v = 2gh for h

lilianamendez2:

do u know how to solve it

callmeizzy1234:

i dont know what you asking

callmeizzy1234:

you can't rearrange the problem to find h because there nothing there to + or x by

jhonyy9:

so you need isolate the term of h on a part for this just divide both sides by 2g and pls. post what will get

jhonyy9:

v = 2gh divide both sides by 2g and get v/2g = (2gh)/2g on the left side what can you make with 2g on numerator and on denominator ?

jhonyy9:

so who posted the problem is offline

callmeizzy1234:

so now you have the problem v/2g=(2gh)/2g

callmeizzy1234:

so now you can combine like terms

jhonyy9:

no pls re-read what i ve wrote above

callmeizzy1234:

okay

callmeizzy1234:

sry

callmeizzy1234:

im stressed

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @callmeizzy1234 so now you have the problem v/2g=(2gh)/2g \(\color{#0cbb34}{\text{End of Quote}}\) what can you make here on the right side ?

jhonyy9:

2g on numerator with 2g on denominator what will do ?

jhonyy9:

@supie